OpenStudy (anonymous):

The product of two consecutive integers is 12 more than the square of the smaller one. Find the integers.

OpenStudy (anonymous):

Any thoughts?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

x = one integer x+1 = second integer Can you work from there?

OpenStudy (johnweldon1993):

\[\large n \times (n + 1) = n^2 + 12\] And can you simplify from that?

OpenStudy (anonymous):

Can you finish it off, with the information we gave you @courtneylvskris ?